OpenStudy (anonymous):

Use the formula for computing future value using compound interest to determine the value of an account at the end of 5 years if a principal amount of $2,500 is deposited in an account at an annual interest rate of 6% and the interest is compounded daily. Assume there are 365 days in a year. Do not round until the end of your calculations; then, round to the nearest cent as needed.

OpenStudy (phoenixfire):

Do you know the compound interest formula?

OpenStudy (tanner23456):

A =P(1+r/n)^nt r = interest rate n= number of times compounded per year t= number of years A=(2500)(1+(.06)/365)^((365)(5)) A=$3,374.56